A racing car has a uniform acceleration of 4 m/s[sup]2[/sup]. What distance will it cover in 10 s after start?
1
Expert's answer
2011-06-09T07:50:35-0400
The distance is obtained by using the following formula:
a=4 m/s2, t=10 s

x=a* t2 / 2=4*102 / 2=200 m.